OpenStudy (anonymous):

What is the volume of a pyramid with base area B = 6 ft^2 and height h = 13 ft? A. 26 ft^3 B. 39 ft^3 C. 52 ft^3 D. 78 ft^3 @DEEISME

OpenStudy (cwrw238):

area = 1/3 * height * area of base

OpenStudy (anonymous):

is it square based pyramid

OpenStudy (anonymous):

idk

OpenStudy (cwrw238):

the shape of the base does not matter . The formula holds for any shaped base

OpenStudy (anonymous):

ok thanks i wasnt know it :)

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

i still don't know the answer

OpenStudy (cwrw238):

well just plug in the values into the formula volume = 1/3 * h * a height h = 13 area a = 6 so its 1/3 * 13 * 6

OpenStudy (anonymous):

are you sure it's not 1/2 * 13 * 6?

OpenStudy (cwrw238):

yes certain - google it if you like...